OpenStudy (anonymous):

Could someone please check my answers ( questions and answers in following post)

OpenStudy (anonymous):

1) Outdoor steps generally require larger landings and lower pitches. A particular outdoor staircase climbs a total vertical elevation of 13 feet from the road level to the front entrance of a building. If the angle of elevation is to be 25 degrees and the landing of each step is to be 28 inches, calculate the following (you must show all work to receive full credit): a) What is the vertical rise per step? (In inches rounded to the nearest whole inch) tan 25°= a/ 28 a/28= .466307658 a/28 * 28= .466307658 * 28 a = (rounded) 13 inches b) Find the number of steps needed to climb from road level to the front entrance. Rise= 13 ft. 13ft=156 in. 156 in./13 in. (Step Height)= 12, 12 steps

OpenStudy (kohai):

 First part is right, though I have no idea how you got there cx